OpenStudy (anonymous):

Why is \[\frac{dy}{dx}=y\]an exact ODE? When I do the math, I obtain this: \[\frac{dy}{dx}=y,\]\[dy=ydx,\]\[-ydx+dy=0,\]This, according to its definition, is not an exact ODE because \[\frac{∂}{∂y}[-y]=-1≠0=\frac{∂}{∂x}[1].\]My professor claims it is exact, but I do not understand why.

myininaya (myininaya):

i don't see it as exact either

myininaya (myininaya):

the test didn't work just like you said

myininaya (myininaya):

oh wait maybe it is depends on how you write it

myininaya (myininaya):

\[\frac{1}{y} \frac{dy}{dx}=1\] \[-1+\frac{1}{y} \frac{dy}{dx}=0\]

myininaya (myininaya):

\[M=-1=>M_y=0\] \[N=\frac{-1}{y}=> N_x=0\]

myininaya (myininaya):

since \[M_y=N_x\], it is exact

OpenStudy (anonymous):

I didnt get exact too. \[M_{y}=-1 \] \[N_{x}=1/y\] *im curious of my answer too.:P

myininaya (myininaya):

but we do get it is exact arranging it like i did

myininaya (myininaya):

we can write it as -1+y'/y=0 like i did above try your test now

OpenStudy (anonymous):

But isn't \[\frac{\partial }{\partial x}\left [ \frac{1}{y} \right ]=\frac{\partial }{\partial x}\left [ y^{-1} \right ]=-y^{-2}y'\]because y is a function of x?

myininaya (myininaya):

remember we treat y like a constant if we are taking the partial with respect to x

myininaya (myininaya):

\[N=\frac{-1}{y} => N_x=0 \]

OpenStudy (anonymous):

Last question, how can you do this: \[\frac{dy}{dx}=y,\]\[\frac{1}{y}\frac{dy}{dx}=1?\]What if \[y=0?\]

myininaya (myininaya):

im assuming y does not equal zero